A monopolists who owns a music streaming service must decide on a price for its yearly subscription. He knows that the demand for music streaming subscription is ??(??)=400−2??PA(QA)=400−2QA for adults and ??(??)=300−??2PT(QT)=300−QT2 for teenagers. The marginal cost of providing music is $10$10

a) (8) Find the profit-maximizing monopoly pricing if they cannot differentiate between teenagers and adults.

b) (7) Find the profit-maximizing monopoly pricing if they can give a discount to teenagers.
Answer #1

Price Discrimination
PA = 400-2QA                     PT = 300-QT2

Given MC = $10

a) Without price discrimination (P= PA = PT )
thus, P =(700 –2Q -Q2)
MR = (700-4Q-3Q2)

For profit maximization MR = MC
                      (700-4Q-3Q2)=10
                      3Q2+4Q-690 = 0
Q = 14.51 thus the optimum level of production is 14.51 unit
                                        P = 700 –2(14.51) –(14.512) =$460.44

b) With price discrimination (PA is not equal to PT)
Thus for adults,
PA = 400-2QA                     
MRA= 400-4QA
For profit maximization MRA = MC
                               400-4QA = 10
required QA= 97.5 thus the optimum level of production for adult is 97.5 unit
   required
PA = (400-(2*97.5)) =$205

Thus for teenagers,
PT = 300-QT2
MRT= 300-3 QT2
For profit maximization MRT = MC
  300-3 QT2=10
                required QT= 9.83 thus the optimum level of production for teenagers is 9.83 unit
                required
PT= (300-9.832)= $203.37